A 0.300 kg potato is tied to a string with length 2.30 m , and the other end of the string is tied to a rigid support. The potato is held straight out horizontally from the point of support, with the string pulled taut, and is then released.

Part A) What is the speed of the potato at the lowest point of its motion? let g be = 9.80 m/s^2

Part B) What is the tension on the string at this point?
Answer #1

Ag given in the question,

Mass of potato: m = 0.3 kg

Radius of circle: r = 2.3 m

(A) Equating potential energy at the start to kinetic energy at the lowest point:

m*g*r = m*v^2 / 2

=>   v^2 = 2*g*r

=> v = sqrt(2*g*r) = sqrt(2*9.8*2.3) = 6.71 m/s

(B) Now, equating the forces to get the tension required:

T - m*g = mv^2 / r

=> T = m*{g + (v^2 / r)} = 0.3*{9.8 + (6.71^2 / 2.3)} = 8.81 N